If y varies directly as x^2, and y = 2 when x = 1, find y when x = 10
Answer by Alan3354(69443)

If y varies directly as x^2, and y = 2 when x = 1, find y when x = 10
-----------
y = kx^2
2 = k*1 --> k = 2
-----
y = 2*10^2 = 200